# Rapid-stable Collective Charging and Reverse Suppressing in Many-Body Quantum Batteries

### v1

@@ I Introduction
The rapid development of quantum technologyL. et al. (2016); Xu et al. (2020) has driven the miniaturization of electronic devices, inspiring many innovations in this field. Quantum batteries (QBs)Alicki and Fannes (2012); Andolina et al. (2018a), as atomic-sized energy storage and conversion devices, are a typical example of such innovations. The concept was first introduced by Robert Alicki and his coauthorsAlicki and Fannes (2012) in 2013. Researchers attempt to reveal whether quantum properties can accelerate the charging processRossini et al. (2020a); Ghosh2021FastCO; Santos (2021), and two charging schemes were proposed: a parallel charging scheme, where each subsystem is independently manipulatedGhosh2021FastCO; Santos (2021), and a collective charging scheme, where global unitary operations are applied to the entire Hilbert space of all subsystemsRossini et al. (2020b); Gyhm et al. (2021); Ghosh et al. (2021). And the ultimate goal of all these work is to leverage quantum effects to enhance battery charging processesCampaioli et al. (2018); Farina et al. (2019); Rossini et al. (2019); Julià-Farré et al. (2020), potentially driving technological advancements across various sectors. Such as, Ref.Tirone et al. (2023) investigates work extraction processes from noisy QBs, focusing on the role of nonlocal resources in enhancing efficiency. The results indicate that while noise generally degrades system performance, properly utilizing nonlocal quantum correlations can mitigate its effects and optimize work extraction. An et al.Song et al. (2023) explored the concept of remote charging for QBs, aiming to enhance their performance by addressing the issue of degradation during the charging process, and the authors propose methods to suppress such degradation, thus improving the overall efficiency and longevity of QBs, with potential applications in quantum technologies. A collision model for the charging of a QB was presented by identical nonequilibrium qubit unitsSeah et al. (2021). It shows that when qubits exhibit quantum coherence, interference effects accelerate the energy distribution spread, leading to faster charging compared to classical incoherent protocols. The study also demonstrates that coherent strategies can achieve higher charging power and characterizes the extractable work using the concept of ergotropy.
@@ IV Charge-discharge dynamics of collective three-qubit QBs
$\hat{H}_{s}=\begin{bmatrix}\varepsilon_{1}&\Omega_{12}(t)&\Omega_{13}(t)&\Omega_{14}(t)\\ \Omega_{21}(t)&\varepsilon_{2}&T_{e}&0\\ \Omega_{31}(t)&T_{e}&\varepsilon_{3}&T_{e}\\ 0&0&T_{e}&\varepsilon_{4}\end{bmatrix}$ (9)
@@ V.1 Possible experimental realization
QBs are energy-efficient storage devices that operate based on quantum tunneling effectsMondal and Bhattacharjee (2021), quantum entanglementImai et al. (2023), and superconducting materialsDou and Yang (2023). Charging refers to the process of transitioning quantum bits from a low-energy state to a high-energy state, while discharging is the reverse process. They can take various physical forms, such as ions, neutral atoms, or photons. In this work, we have proposed a many-body QB model with a tunable energy gap. Theoretically, any quantum devices with the above-mentioned energy gap could serve as an experimental candidate for this model. However, there is a prerequisite that should be met: the gap and tunneling effect parameters of the individual QBs mentioned are the same. Therefore, the physical materials of the individual QBs forming the many-body QB need to be identical in order to achieve the theoretical results predicted by this model. Both two-level systems with appropriate band gaps and quantum spin systems can serve as potential candidates for the experimental realization of this work.

### v2

@@ I Introduction
The rapid development of quantum technologyL. et al. (2016); Xu et al. (2020) has driven the miniaturization of electronic devices, inspiring many innovations in this field. Quantum batteries (QBs)Alicki and Fannes (2012); Andolina et al. (2018a), as atomic-sized energy storage and conversion devices, are a typical example of such innovations. The concept was first introduced by Robert Alicki and his coauthorsAlicki and Fannes (2012) in 2013. Researchers attempt to reveal whether quantum properties can accelerate the charging processRossini et al. (2020a); Santos (2021), and two charging schemes were proposed: a parallel charging scheme, where each subsystem is independently manipulatedSantos (2021), and a collective charging scheme, where global unitary operations are applied to the entire Hilbert space of all subsystemsRossini et al. (2020b); Gyhm et al. (2021). And the ultimate goal of all these work is to leverage quantum effects to enhance battery charging processesCampaioli et al. (2018); Farina et al. (2019); Rossini et al. (2019); Julià-Farré et al. (2020), potentially driving technological advancements across various sectors. Such as, Ref.Tirone et al. (2023) investigates work extraction processes from noisy QBs, focusing on the role of nonlocal resources in enhancing efficiency. The results indicate that while noise generally degrades system performance, properly utilizing nonlocal quantum correlations can mitigate its effects and optimize work extraction. An et al.Song et al. (2023) explored the concept of remote charging for QBs, aiming to enhance their performance by addressing the issue of degradation during the charging process, and the authors propose methods to suppress such degradation, thus improving the overall efficiency and longevity of QBs, with potential applications in quantum technologies. A collision model for the charging of a QB was presented by identical nonequilibrium qubit unitsSeah et al. (2021). It shows that when qubits exhibit quantum coherence, interference effects accelerate the energy distribution spread, leading to faster charging compared to classical incoherent protocols. The study also demonstrates that coherent strategies can achieve higher charging power and characterizes the extractable work using the concept of ergotropy.
@@ IV Charge-discharge dynamics of collective three-qubit QBs
$\hat{H}_{s}=\begin{bmatrix}\varepsilon_{1}&\Omega_{12}(t)&\Omega_{13}(t)&\Omega_{14}(t)\\ \Omega_{21}(t)&\varepsilon_{2}&T_{e}&0\\ \Omega_{31}(t)&T_{e}&\varepsilon_{3}&T_{e}\\ \Omega_{41}(t)&0&T_{e}&\varepsilon_{4}\end{bmatrix}$ (9)
@@ V.1 Possible experimental realization
QBs are energy-efficient storage devices that operate based on quantum tunneling effectsMondal and Bhattacharjee (2021), quantum entanglementShi et al. (2022), and superconducting materialsDou and Yang (2023). Charging refers to the process of transitioning quantum bits from a low-energy state to a high-energy state, while discharging is the reverse process. They can take various physical forms, such as ions, neutral atoms, or photons. In this work, we have proposed a many-body QB model with a tunable energy gap. Theoretically, any quantum devices with the above-mentioned energy gap could serve as an experimental candidate for this model. However, there is a prerequisite that should be met: the gap and tunneling effect parameters of the individual QBs mentioned are the same. Therefore, the physical materials of the individual QBs forming the many-body QB need to be identical in order to achieve the theoretical results predicted by this model. Both two-level systems with appropriate band gaps and quantum spin systems can serve as potential candidates for the experimental realization of this work.